# ChatCFD: An LLM-Driven Agent for End-to-End CFD Automation with Domain-Specific Structured Reasoning

### v2

@@ ACKNOWLEDGEMENTS
E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft. Kang Hu: Methodology, Software, Data curation, Formal analysis, Resources. Zhuowen Wu: Methodology. Jiangyang Ge: Data curation, Formal analysis. Jiawei Miao: Software, Resources. Yuzhi Zhang: Conceptualization, Software, Resources. He Sun: Resources, Writing - Review. Weizong Wang: Resources, Writing - Review, Funding acquisition. Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
@@ Appendix B Appendix: CFD results of Cases 1 to 5
Figures 12 and 13 show the velocity magnitude contours and surface pressure coefficient distributions for the NACA0012 airfoil Cases 1 to 4 at 10° angle of attack using different turbulence models. As can be seen, the velocity contours from the four cases are similar, and the pressure coefficient distributions agree well with the experimental results [37].

### v3

@@ ACKNOWLEDGEMENTS
E Fan: Methodology, Data curation, Formal analysis, Writing - original draft. Kang Hu: Methodology, Software, Data curation, Formal analysis, Writing - original draft. Zhuowen Wu: Methodology. Jiangyang Ge: Data curation, Formal analysis. Jiawei Miao: Software, Resources. Yuzhi Zhang: Conceptualization, Software, Resources. He Sun: Resources, Writing - Review. Weizong Wang: Resources, Writing - Review, Funding acquisition. Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
@@ Appendix B Appendix: CFD results of Cases 1 to 5
Figures 17 and 18 show the velocity magnitude contours and surface pressure coefficient distributions for the NACA0012 airfoil Cases 1 to 4 at 10° angle of attack using different turbulence models. As can be seen, the velocity contours from the four cases are similar, and the pressure coefficient distributions agree well with the experimental results [11].